OpenStudy (anonymous):

what equation in slop intercept form represents the line that passes through (1,-7) and (3,1)? please help

OpenStudy (freckles):

well let's begin by finding the slope

OpenStudy (freckles):

do you know how to find the slope of a line given two points on that line?

OpenStudy (anonymous):

y=mx+b?

OpenStudy (freckles):

we will use y=mx+b to setup the equation but we want to find the values m and b. we are trying to find m right now m is known as the slope of this equation you can find the slope by doing the following : If given two points on a line let's call them (a,b) and (c,d) then the slope of that line is: \[\frac{d-b}{c-a}\]

OpenStudy (freckles):

you may have seen this formula written with different letters like maybe you seen it as \[\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (freckles):

you can line up the points if you want to like this: (1,-7) (3,1) --- and subtract but just remember to put your y difference over your x difference to find the slope

OpenStudy (freckles):

so -7-1=? and 1-3=?

OpenStudy (freckles):

right -8/-2 is 8/2

OpenStudy (freckles):

you can reduce 8/2 though

OpenStudy (freckles):

we know this can be reduced since we know that 2 divides 8 evenly

OpenStudy (freckles):

so we are going to replace m in y=mx+b with?

OpenStudy (freckles):

why did you delete your comment ? :p